# Venus: An Efficient Edge Memory-and-Retrieval System for VLM-based Online Video Understanding

### v1

@@ I Introduction
Building on the advancements of large language models, vision-language models (VLMs) [liu2023visual, zhang2024video] have been pre-trained on vast collections of image-text pairs, allowing them to learn rich representations that capture fine-grained visual-linguistic correlations. These capabilities have driven significant progress across various visual understanding tasks [tang2025adaptive, feng2025vision]. With these capabilities, an increasing number of commercial and research organizations are self-hosting VLMs on public or private cloud infrastructures and exposing them via standardized APIs [openai_api]. This growing accessibility to advanced vision-language reasoning has empowered developers and researchers to integrate VLMs into real-world systems, fueling the rapid development of edge intelligence applications that advance visual understanding, including multimodal personal assistants [li2024personal], smart surveillance [fan2025ai], and city scene reasoning [sun20243d].
Edge cameras are ubiquitous, spanning a wide range of public and private infrastructures. Online video from edge cameras is continuously recorded and retrieved on demand, making intelligent analysis and understanding of streaming edge video a longstanding focus in edge AI research [hsieh2018focus, xu2020approximate]. Prior to the VLM era, CNNs had long served as the backbone of online video understanding. While lightweight, CNNs are insufficient for high-level semantic retrieval and cross-modal reasoning, driving a paradigm shift toward VLM-based workflows that incorporate rich visual-linguistic reasoning [niu2025ovo, huang2025online]. Figure 1 abstracts a VLM-powered online video understanding application deployed in a smart-home scenario, where family members can initiate queries targeting current or historical segments of recorded video at any time. The application consists of three key modules: (1) Streaming Perception Module: Edge cameras equipped with onboard compute continuously capture video of household scenes and process multimodal data on the fly. (2) Historical Memory Module: Historical video streams are constructed and stored in a memory database to support efficient and accurate retrieval of query-relevant information. (3) Reasoning Module: Upon receiving a user query, cloud-hosted VLMs are leveraged to perform reasoning and generate answers, with historical memory providing contextual grounding.
@@
Emerging research efforts in the AI community [liu2025bolt, li2024llava, luo2024video] have explored the effectiveness of VLMs for video understanding, but most efforts focus on algorithmic performance while overlooking deployment-related constraints, limiting their direct applicability in edge intelligence scenarios. Fully cloud-based deployment leverages powerful server-side computation to reduce on-device processing, but requires uploading the entire relevant video, incurring significant communication latency. Some methods [liu2025bolt, tang2025adaptive] improve reasoning accuracy by performing frame selection prior to inference. An alternative deployment strategy executes the selection algorithm on the edge, allowing only the selected frames to be uploaded to the cloud. While this significantly reduces communication overhead, the limited compute capacity of edge devices incurs substantial on-device processing costs. Beyond overwhelmed communication and computation overhead, existing methods are inherently restricted to offline execution and are unsuitable for online streaming applications that involve online user queries and require long-term contextual memory.
@@ II-A VLM-Based Online Video Understanding Applications
VLMs demonstrate strong capabilities in multimodal perception [liu2023visual, zhang2024video] and are increasingly integrated into edge applications to enhance visual understanding. A VLM-based online video understanding application employs an edge camera for continuous video capture, enabling natural language queries over current or historical video segments at any time, with real-time, accurate responses to users powered by a VLM. For example, queries may involve recalling recent activities such as the cooking process or verifying whether an elderly person took their medication in the afternoon.
Difference with Traditional Video Analysis Workflow. Prior to the VLM era, CNNs had long served as the backbone of traditional online video analysis systems. Numerous on-device [li2020reducto, hsieh2018focus] or edge–cloud collaborative [wang2024gecko, zhang2021towards, yan2024visflow] studies have proposed specialized optimizations for CNN-based real-time video analytics pipelines. However, fundamental architectural and functional disparities between CNNs and VLMs hinder the transferability of optimizations from CNN-based to VLM-based pipelines. Also, VLMs remain impractical for edge deployment due to their substantially higher resource demands than CNNs, and many state-of-the-art VLMs are still closed-source, making them hard to self-host. These factors result in VLM capabilities being typically accessible only via black-box APIs hosted by cloud providers. These differences motivate the redesign of workflows for the VLM era.
@@ II-B Limitations of Existing VLM-based Video Understanding
Recent advances in the AI community [li2024llava, zhang2023video] have introduced VLMs into video understanding tasks to enhance semantic representation and cross-modal integration. Existing VLM-based video understanding methods generally overlook deployment-related constraints, which prevent their direct applicability in edge intelligence scenarios.
LLaVA-OneVision [li2024llava] processes every video frame using 196 tokens without any reduction, resulting in approximately 6 million tokens for a one-hour video at 8 FPS. This approach introduces significant video transmission and inference overhead. Video-RAG [luo2024video] and MDF [han2024self] apply uniform sampling or redundant frame removal to reduce video token redundancy. Although efficient, these query-agnostic strategies assume equal frame importance and often underperform in complex real-world scenarios. Sampling too sparsely risks missing critical content, while oversampling introduces redundancy that increases transmission and inference costs.
Other recent methods such as BOLT [liu2025bolt] and AKS [tang2025adaptive] leverage vision-language similarity to identify frames relevant to the input query. They typically use contrastive language-image pretraining encoders [radford2021learning] to associate frames with textual descriptions, followed by tailored selection algorithms that only retrieve query-relevant frames for downstream video understanding with VLMs. Their adaptive selection helps balance informativeness and inference efficiency in video understanding. However, these methods still require transferring the entire relevant video to the cloud for processing, making significant communication latency unavoidable. An alternative is to execute these algorithms on edge devices, allowing only the selected frames to be uploaded to the cloud, which can significantly reduce transmission overhead. However, such a frame-wise algorithm design is impractical for online scenarios, as edge devices often lack the computational capacity to run transformer-based encoders on every frame in real time.
@@ III-A1 Multimodal Embedding Models
Embedding was originally introduced to transform data such as text, images, or audio into vector representations, enabling similarity comparisons within the same modality [lin2017structured, dosovitskiy2020image]. However, unimodal embedding is inherently limited in capturing cross-modal information. To bridge this gap, multimodal embedding models (MEMs) [zhou2024megapairs, radford2021learning, zhai2023sigmoid] have been developed to map different modalities into a shared representation space. By optimizing the alignment of representations across modalities [^1] This work focuses on visual and linguistic modalities, specifically image and text, for video understanding. Although other modalities (e.g., audio, temperature signal) are not included, Venus can be naturally extended by integrating MEMs that support additional modalities, such as ImageBind [girdhar2023imagebind]., MEMs facilitate semantic interaction between modalities through a shared embedding space. In the context of online video understanding, this enables user-friendly querying via natural language to efficiently retrieve query-relevant video frames.
@@ III-A2 Vector Database
MEMs enable embedding multimodal data into a unified vector space, where vector databases enable efficient storage and fast nearest neighbor retrieval for similarity search [wang2025timestamp]. Given a set of data vectors $\{x_{i}\}$ and a query $q$, the goal is to identify the most similar vectors to $q$, where similarity is typically defined using either Euclidean distance or inner-product. Facebook AI Similarity Search (FAISS) [johnson2019billion] exemplifies such systems, accelerating nearest neighbor search via clustering, quantization, graph-based indexing, and hardware acceleration for scalable, low-latency retrieval in large vector spaces.
@@ III-C Technical Challenges in Practical Deployment
High latency hinders real-time embedding and ingestion. Typical edge cameras capture video at 25–60 FPS [TL_IPC632P_A4], which imposes substantial overhead when applying MEMs for frame-wise embedding, particularly on resource-constrained edge devices. When the input FPS exceeds the device’s embedding throughput, frames accumulate in a backlog. Upon receiving a user query at timestamp $t$, the system must first embed all pending frames before $t$ before performing retrieval, introducing latency and degrading responsiveness.
To assess the feasibility of real-time on-device embedding, we evaluate three representative edge platforms: NVIDIA Jetson AGX Orin [jetson-orin], Jetson Xavier NX [jetson-xavier], and Jetson TX2 [jetson-tx2] under realistic streaming conditions, where frames are processed on the fly as they are captured. As shown in Figure 4, embedding latency increases rapidly with higher FPS, making real-time embedding infeasible beyond a certain FPS threshold. At the native frame rate of typical network cameras (e.g., 25 FPS), embedding delay exceeds 212 minutes, which is clearly impractical for online query. Our results indicate that real-time embedding and ingestion are only achievable when the frame rate is drastically reduced, such as 0.3 FPS for Jetson TX2, 0.7 FPS for Jetson Xavier NX, and 1.8 FPS for Jetson AGX Orin. However, such aggressive uniform downsampling leads to severe information loss. If key frames required to answer user queries are discarded, the system may entirely fail to perform accurate video understanding.
@@ IV-B Scene Segmentation and Clustering for Frame Filtering
Streaming video at 25–60 FPS [TL_IPC632P_A4] on typical edge cameras makes frame-wise perception and memory prohibitively expensive for real-time processing. Uniform downsampling is widely used in prior work, but often results in severe and uncontrollable information loss. Venus aims not to prune frames aggressively, but instead to preserve raw video and construct a sparse index efficiently using lightweight online algorithms. The index comprises representative frames that summarize visually similar content, enabling efficient localization of query-relevant scenes and subsequent fine-grained sampling during the querying stage. This section focuses on selecting indexed frames through a scene segmentation and frame clustering module, which consists of two steps: scene segmentation and frame clustering, as illustrated in Figure 7.
@@ IV-B1 Scene Detection and Segmentation
Venus takes streaming video as input and incrementally partitions it into contiguous temporal segments using a coarse-grained algorithm, termed scene detection and segmentation. We perform a lightweight coarse-grained partitioning of the video stream using visual features including hue, saturation, brightness, and edge maps [cao2006density]. Specifically, we define a frame-level metric $\phi$ to quantify visual differences between consecutive frames based on intrinsic pixel-level features. Given a streaming video $\mathcal{F}=\{f_{0},f_{1},\dots\}$, where $f_{i}$ denotes the frame at time step $i$, the scene tracking score for $f_{i}$ is computed as:
@@ IV-B2 Frame Clustering
We perform frame clustering within each scene partition for two main reasons: First, grouping visually similar frames at a finer granularity further reduces redundancy. Second, cluster centroids serve as representative frames that summarize similar content, making them natural candidates for building a sparse index. Existing research [tang2025adaptive] on video analysis also adopts clustering algorithms to reduce redundancy, using methods such as K-Means, DBSCAN, and their variants to group similar frames. However, a limitation of such clustering algorithms is that frames within a cluster may be temporally disjoint, potentially hindering temporal reasoning. To this end, Venus adopts a straightforward incremental clustering strategy, which has been well-studied in the literature [cao2006density, o2002streaming]. Specifically, for each scene partition $m_{i}\in\mathcal{M}$, clustering proceeds as follows: the first frame is assigned to the initial cluster $c_{1}$. To cluster a new frame $f_{i}$, we first flatten its raw pixel values into a vector representation. We then compute the $L_{2}$ distance between this vector and the centroid of each existing cluster. If the minimum distance is within a threshold, $f_{i}$ is assigned to the nearest cluster; otherwise, a new cluster is created with $f_{i}$ as its centroid. After clustering, we obtain a set of clusters $\mathcal{C}=\{c_{0},c_{1},$...\}, from which the centroid frame of each cluster is selected as an indexed frame. These selected indexed frames, being highly sparse compared to the original video stream, are used for embedding and memory index construction, enabling real-time on-device perception.
@@ IV-C1 Memory Construction with MEMs
$\mathrm{AuxModels}$ is a collection of lightweight proprietary models and algorithms, such as OCR [easyocr] and YOLO [redmon2016you], which can be dynamically configured according to application needs and edge device constraints. Their outputs are formatted into predefined textual templates to construct auxiliary prompts. Finally, the indexed frames and their associated auxiliary prompts are fed into the MEMs to generate vector representations within a unified image-text embedding space:
@@ V IMPLEMENTATION AND EVALUATION
We have fully implemented the prototype system of Venus and baselines with $\sim$1000 LoC in Python atop PyTorch. This section presents a comprehensive evaluation of the Venus prototype, including comparisons with baselines across various datasets, models, and edge platforms.
@@ V-A1 Hardware Setup
On the edge side, we used NVIDIA Jetson AGX Orin [jetson-orin] devices for our experiments with each device equipped with an NVMe SSD for external storage. On the cloud side, we use a server with NVIDIA L40S GPUs (48 GB memory) to emulate cloud compute resources, where the VLMs are deployed. The network bandwidth between the edge and cloud is fixed at 100 Mbps, which is typical for wired or wireless edge networks.
@@
Model
Method
Video-MME (Short) Video-MME (Medium) Video-MME (Long) EgoSchema (Full)
@@ V-A2 Datasets and Models
We evaluate Venus on two widely used VQA benchmarks, Video-MME [fu2025video] and EgoSchema [mangalam2023egoschema], which simulate realistic online video understanding scenarios. Video-MME consists of 900 videos and 2,700 questions, with some video clips exceeding one hour in length. EgoSchema contains over 5,000 multiple-choice QA pairs grounded in approximately 250 hours of egocentric video footage. All input videos are uniformly sampled at 8 FPS in our experiments. We assume that the query arrives immediately after the playback of each video clip, simulating an online video understanding scenario in which users query specific temporal segments of the video in real time. To evaluate the performance of VLM-based video understanding, we deploy two popular open-source VLMs on the cloud server: LLaVA-OV-7B [li2024llava] and Qwen2-VL-7B [wang2024qwen2]. We employ 7B models due to hardware constraints, while Venus’s modules are independent of the VLM and invoke it only via API calls for final reasoning, enabling access to cloud-hosted models of any size. We use BGE-VL-large [zhou2024megapairs] as the MEM. We use EasyOCR [easyocr] and YOLO [redmon2016you] as auxiliary models.

### v2

@@ I Introduction
Building on the advancements of large language models, vision-language models (VLMs) [21, 45] have been pre-trained on vast collections of image-text pairs, allowing them to learn rich representations that capture fine-grained visual-linguistic correlations. These capabilities have driven significant progress across various visual understanding tasks [34, 5]. With these capabilities, an increasing number of commercial and research organizations are self-hosting VLMs on public or private cloud infrastructures and exposing them via standardized APIs [29]. This growing accessibility to advanced vision-language reasoning has empowered developers and researchers to integrate VLMs into real-world systems, fueling the rapid development of edge intelligence applications that advance visual understanding, including multimodal personal assistants [18, 41], smart surveillance [4], and city scene reasoning [33].
Edge cameras are ubiquitous, spanning a wide range of public and private infrastructures. Online video from edge cameras is continuously recorded and retrieved on demand, making intelligent analysis and understanding of streaming edge video a longstanding focus in edge AI research [10, 39]. Prior to the VLM era, CNNs had long served as the backbone of online video understanding. While lightweight, CNNs are insufficient for high-level semantic retrieval and cross-modal reasoning, driving a paradigm shift toward VLM-based workflows that incorporate rich visual-linguistic reasoning [27, 11]. Figure 1 abstracts a VLM-powered online video understanding application deployed in a smart-home scenario, where family members can initiate queries targeting current or historical segments of recorded video at any time. The application consists of three key modules: (1) Streaming Perception Module: Edge cameras equipped with onboard compute continuously capture video of household scenes and process multimodal data on the fly. (2) Historical Memory Module: Historical video streams are constructed and stored in a memory database to support efficient and accurate retrieval of query-relevant information. (3) Reasoning Module: Upon receiving a user query, cloud-hosted VLMs are leveraged to perform reasoning and generate answers, with historical memory providing contextual grounding.
@@
Emerging research efforts in the AI community [23, 17, 24] have explored the effectiveness of VLMs for video understanding, but most efforts focus on algorithmic performance while overlooking deployment-related constraints, limiting their direct applicability in edge intelligence scenarios. Fully cloud-based deployment leverages powerful server-side computation to reduce on-device processing, but requires uploading the entire relevant video, incurring significant communication latency. Some methods [23, 34] improve reasoning accuracy by performing frame selection prior to inference. An alternative deployment strategy executes the selection algorithm on the edge, allowing only the selected frames to be uploaded to the cloud. While this significantly reduces communication overhead, the limited compute capacity of edge devices incurs substantial on-device processing costs. Beyond overwhelmed communication and computation overhead, existing methods are inherently restricted to offline execution and are unsuitable for online streaming applications that involve online user queries and require long-term contextual memory.
@@ II-A VLM-Based Online Video Understanding Applications
VLMs demonstrate strong capabilities in multimodal perception [21, 45] and are increasingly integrated into edge applications to enhance visual understanding. A VLM-based online video understanding application employs an edge camera for continuous video capture, enabling natural language queries over current or historical video segments at any time, with real-time, accurate responses to users powered by a VLM. For example, queries may involve recalling recent activities such as the cooking process or verifying whether an elderly person took their medication in the afternoon.
Difference with Traditional Video Analysis Workflow. Prior to the VLM era, CNNs had long served as the backbone of traditional online video analysis systems. Numerous on-device [19, 10] or edge–cloud collaborative [36, 44, 40] studies have proposed specialized optimizations for CNN-based real-time video analytics pipelines. However, fundamental architectural and functional disparities between CNNs and VLMs hinder the transferability of optimizations from CNN-based to VLM-based pipelines. Also, VLMs remain impractical for edge deployment due to their substantially higher resource demands than CNNs, and many state-of-the-art VLMs are still closed-source, making them hard to self-host. These factors result in VLM capabilities being typically accessible only via black-box APIs hosted by cloud providers. These differences motivate the redesign of workflows for the VLM era.
@@ II-B Limitations of Existing VLM-based Video Understanding
Recent advances in the AI community [17, 43] have introduced VLMs into video understanding tasks to enhance semantic representation and cross-modal integration. Existing VLM-based video understanding methods generally overlook deployment-related constraints, which prevent their direct applicability in edge intelligence scenarios.
LLaVA-OneVision [17] processes every video frame using 196 tokens without any reduction, resulting in approximately 6 million tokens for a one-hour video at 8 FPS. This approach introduces significant video transmission and inference overhead. Video-RAG [24] and MDF [9] apply uniform sampling or redundant frame removal to reduce video token redundancy. Although efficient, these query-agnostic strategies assume equal frame importance and often underperform in complex real-world scenarios. Sampling too sparsely risks missing critical content, while oversampling introduces redundancy that increases transmission and inference costs.
Other recent methods such as BOLT [23] and AKS [34] leverage vision-language similarity to identify frames relevant to the input query. They typically use contrastive language-image pretraining encoders [30] to associate frames with textual descriptions, followed by tailored selection algorithms that only retrieve query-relevant frames for downstream video understanding with VLMs. Their adaptive selection helps balance informativeness and inference efficiency in video understanding. However, these methods still require transferring the entire relevant video to the cloud for processing, making significant communication latency unavoidable. An alternative is to execute these algorithms on edge devices, allowing only the selected frames to be uploaded to the cloud, which can significantly reduce transmission overhead. However, such a frame-wise algorithm design is impractical for online scenarios, as edge devices often lack the computational capacity to run transformer-based encoders on every frame in real time.
@@ III-A1 Multimodal Embedding Models
Embedding was originally introduced to transform data such as text, images, or audio into vector representations, enabling similarity comparisons within the same modality [20, 3]. However, unimodal embedding is inherently limited in capturing cross-modal information. To bridge this gap, multimodal embedding models (MEMs) [46, 30, 42] have been developed to map different modalities into a shared representation space. By optimizing the alignment of representations across modalities [^1] This work focuses on visual and linguistic modalities, specifically image and text, for video understanding. Although other modalities (e.g., audio, temperature signal) are not included, Venus can be naturally extended by integrating MEMs that support additional modalities, such as ImageBind [8]., MEMs facilitate semantic interaction between modalities through a shared embedding space. In the context of online video understanding, this enables user-friendly querying via natural language to efficiently retrieve query-relevant video frames.
@@ III-A2 Vector Database
MEMs enable embedding multimodal data into a unified vector space, where vector databases enable efficient storage and fast nearest neighbor retrieval for similarity search [38]. Given a set of data vectors $\{x_{i}\}$ and a query $q$, the goal is to identify the most similar vectors to $q$, where similarity is typically defined using either Euclidean distance or inner-product. Facebook AI Similarity Search (FAISS) [16] exemplifies such systems, accelerating nearest neighbor search via clustering, quantization, graph-based indexing, and hardware acceleration for scalable, low-latency retrieval in large vector spaces.
@@ III-C Technical Challenges in Practical Deployment
High latency hinders real-time embedding and ingestion. Typical edge cameras capture video at 25–60 FPS [35], which imposes substantial overhead when applying MEMs for frame-wise embedding, particularly on resource-constrained edge devices. When the input FPS exceeds the device’s embedding throughput, frames accumulate in a backlog. Upon receiving a user query at timestamp $t$, the system must first embed all pending frames before $t$ before performing retrieval, introducing latency and degrading responsiveness.
To assess the feasibility of real-time on-device embedding, we evaluate three representative edge platforms: NVIDIA Jetson AGX Orin [13], Jetson Xavier NX [15], and Jetson TX2 [14] under realistic streaming conditions, where frames are processed on the fly as they are captured. As shown in Figure 4, embedding latency increases rapidly with higher FPS, making real-time embedding infeasible beyond a certain FPS threshold. At the native frame rate of typical network cameras (e.g., 25 FPS), embedding delay exceeds 212 minutes, which is clearly impractical for online query. Our results indicate that real-time embedding and ingestion are only achievable when the frame rate is drastically reduced, such as 0.3 FPS for Jetson TX2, 0.7 FPS for Jetson Xavier NX, and 1.8 FPS for Jetson AGX Orin. However, such aggressive uniform downsampling leads to severe information loss. If key frames required to answer user queries are discarded, the system may entirely fail to perform accurate video understanding.
@@ IV-B Scene Segmentation and Clustering for Frame Filtering
Streaming video at 25–60 FPS [35] on typical edge cameras makes frame-wise perception and memory prohibitively expensive for real-time processing. Uniform downsampling is widely used in prior work, but often results in severe and uncontrollable information loss. Venus aims not to prune frames aggressively, but instead to preserve raw video and construct a sparse index efficiently using lightweight online algorithms. The index comprises representative frames that summarize visually similar content, enabling efficient localization of query-relevant scenes and subsequent fine-grained sampling during the querying stage. This section focuses on selecting indexed frames through a scene segmentation and frame clustering module, which consists of two steps: scene segmentation and frame clustering, as illustrated in Figure 7.
@@ IV-B1 Scene Detection and Segmentation
Venus takes streaming video as input and incrementally partitions it into contiguous temporal segments using a coarse-grained algorithm, termed scene detection and segmentation. We perform a lightweight coarse-grained partitioning of the video stream using visual features including hue, saturation, brightness, and edge maps [2]. Specifically, we define a frame-level metric $\phi$ to quantify visual differences between consecutive frames based on intrinsic pixel-level features. Given a streaming video $\mathcal{F}=\{f_{0},f_{1},\dots\}$, where $f_{i}$ denotes the frame at time step $i$, the scene tracking score for $f_{i}$ is computed as:
@@ IV-B2 Frame Clustering
We perform frame clustering within each scene partition for two main reasons: First, grouping visually similar frames at a finer granularity further reduces redundancy. Second, cluster centroids serve as representative frames that summarize similar content, making them natural candidates for building a sparse index. Existing research [34] on video analysis also adopts clustering algorithms to reduce redundancy, using methods such as K-Means, DBSCAN, and their variants to group similar frames. However, a limitation of such clustering algorithms is that frames within a cluster may be temporally disjoint, potentially hindering temporal reasoning. To this end, Venus adopts a straightforward incremental clustering strategy, which has been well-studied in the literature [2, 28]. Specifically, for each scene partition $m_{i}\in\mathcal{M}$, clustering proceeds as follows: the first frame is assigned to the initial cluster $c_{1}$. To cluster a new frame $f_{i}$, we first flatten its raw pixel values into a vector representation. We then compute the $L_{2}$ distance between this vector and the centroid of each existing cluster. If the minimum distance is within a threshold, $f_{i}$ is assigned to the nearest cluster; otherwise, a new cluster is created with $f_{i}$ as its centroid. After clustering, we obtain a set of clusters $\mathcal{C}=\{c_{0},c_{1},$...\}, from which the centroid frame of each cluster is selected as an indexed frame. These selected indexed frames, being highly sparse compared to the original video stream, are used for embedding and memory index construction, enabling real-time on-device perception.
@@ IV-C1 Memory Construction with MEMs
$\mathrm{AuxModels}$ is a collection of lightweight proprietary models and algorithms, such as OCR [12] and YOLO [31], which can be dynamically configured according to application needs and edge device constraints. Their outputs are formatted into predefined textual templates to construct auxiliary prompts. Finally, the indexed frames and their associated auxiliary prompts are fed into the MEMs to generate vector representations within a unified image-text embedding space:
@@ V-A1 Hardware Setup
On the edge side, we used NVIDIA Jetson AGX Orin [13] devices for our experiments with each device equipped with an NVMe SSD for external storage. On the cloud side, we use a server with NVIDIA L40S GPUs (48 GB memory) to emulate cloud compute resources, where the VLMs are deployed. The network bandwidth between the edge and cloud is fixed at 100 Mbps, which is typical for wired or wireless edge networks.
@@
  Model Method Video-MME (Short) Video-MME (Medium) Video-MME (Long) EgoSchema (Full)
@@ V-A2 Datasets and Models
We evaluate Venus on two widely used VQA benchmarks, Video-MME [6] and EgoSchema [26], which simulate realistic online video understanding scenarios. Video-MME consists of 900 videos and 2,700 questions, with some video clips exceeding one hour in length. EgoSchema contains over 5,000 multiple-choice QA pairs grounded in approximately 250 hours of egocentric video footage. All input videos are uniformly sampled at 8 FPS in our experiments. We assume that the query arrives immediately after the playback of each video clip, simulating an online video understanding scenario in which users query specific temporal segments of the video in real time. To evaluate the performance of VLM-based video understanding, we deploy two popular open-source VLMs on the cloud server: LLaVA-OV-7B [17] and Qwen2-VL-7B [37]. We employ 7B models due to hardware constraints, while Venus’s modules are independent of the VLM and invoke it only via API calls for final reasoning, enabling access to cloud-hosted models of any size. We use BGE-VL-large [46] as the MEM. We use EasyOCR [12] and YOLO [31] as auxiliary models.